OpenStudy (anonymous):

Tarzan is often portrayed as swinging through the jungle from vine to vine. Each of Tarzan's vines are 30 meters long and Tarzan catches each vine when it is at rest 15 degrees to the left of the (downwards) vertical, swings on it until it stops at an angle of 15 degrees to the right of vertical, and then grabs the next vine at rest and repeats the process. What is Tarzan's average horizontal speed in m/s? Details and assumptions You may take g to be 9.8 m/s2. You may treat Tarzan as a simple pendulum and use the small angle approximation.

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

 Consider this one swing. The time it takes to swing from one side to the other is half the time period of the pendulum, right? So, \(\large t=\frac{T}{2}\) Since you can use small angle approximation, you can take 'd' as the distance traveled by tarzan. So, average horizontal velocity=\(\Large \frac{d}{t}\) All you have to do is find 'd' and 't'. Call me if you don't know how to find them.

OpenStudy (anonymous):

i have found d using cosine rule how to find t

OpenStudy (anonymous):

T is given by the formula \[T=2\pi\sqrt(\frac{L}{g})\] 't' then is just \(\large \frac{T}{2}\).